# On the convergence of loss and uncertainty-based active learning algorithms

### v2

@@ 3.1 Bernoulli Sampling with Constant Step Size: First Set of Convergence Conditions
Next, we consider the generalized smooth hinge loss function (Rennie, 2005), defined for $a\geq 1$ as follows:
$\ell(u)=\left\{\begin{array}[]{ll}\frac{a}{a+1}-u&\text{ if }u\leq 0\\ \frac{a}{a+1}-u+\frac{1}{a+1}u^{a+1}&\text{ if }0\leq u\leq 1\\ 0&\text{ otherwise}.\end{array}\right.$ (6)
This is a continuously differentiable function that converges to the value of the hinge loss function, $\max\{1-u,0\}$, as $a$ goes to infinity. The family of loss functions parameterized by $a$ accommodates the smooth hinge loss function with $a=1$, introduced by Rennie and Srebro (2005).
Theorem 3.4.
$\pi^{*}(u)=\left\{\begin{array}[]{ll}\beta\frac{\frac{a}{a+1}-u}{\rho-u}&\text{ if }u\leq 0\\ \beta\frac{\frac{a}{a+1}\frac{1}{1-u^{a}}(1-u)-\frac{1}{a+1}u}{\rho-u}&\text{ if }0\leq u\leq 1\\ 0&\text{ if }u\geq 1.\end{array}\right.$
Note that the sampling probability function $\pi^{*}$ is increasing in $a$ and is upper-bounded by $\pi^{**}(u)=\beta(1-u)/(\rho-1)$, which may be regarded as the limit for the hinge loss function. See Figure 1 for an illustration.
We remark that the expected loss bound in Theorem 3.4 is the same as for the squared hinge loss function in Theorem 3.3, except for an additional factor $c_{a,\rho}=a/(a(\rho-1)+1)$. This factor is increasing in $a$ but always lies in the interval $[1/\rho,1/(\rho-1)]$, where the boundary values are achieved for $a=1$ and $a\rightarrow\infty$, respectively.
Theorem 3.5.
Assuming $\rho^{*}>1$, the loss function is the generalized smooth hinge loss, and the sampling probability function $\pi^{*}$ satisfies the assumptions in Theorem 3.4, with $n>((a+1)/a)c_{a,\rho}R^{2}S^{2}/\beta$, the expected number of sampled points is bounded as
where $\kappa$ is some positive constant.
For any fixed number of iterations $n$ satisfying the condition of the theorem, the expected number of sampled points is bounded by a constant for sufficiently large value of parameter $a$. The bound in Theorem 3.5 depends on how the loss function $\ell(u)$ varies with $u$. When $a$ is large, $\ell(u)$ is approximately $1-u$ (hinge loss), otherwise, it is approximately $\frac{a}{2}(1-u)^{2}$ for $0\leq u\leq 1$ (squared hinge loss).
Multi-class Classification
The conditions for multi-class classification with the set of classes $\mathcal{Y}$ are the same as for binary classification, with $u(x,y,\theta):=x^{\top}\theta_{y}-\max_{y^{\prime}\in\mathcal{Y}\setminus\{y\}}x^{\top}\theta_{y^{\prime}}$, except for an additional factor of $2$ in the left-hand side of the inequality in the first condition (see Lemma A.5 in the Appendix). Hence, all the observations remain valid for the multi-class classification case.
Discussion
The conditions in Assumption 15 allow us to obtain convergence rate bounds for linearly separable data for some other predictors and loss functions. However, for linear classifiers, loss functions must satisfy certain properties related to hinge loss function. Further details are discussed in Appendix A.7. We next present a framework for deriving convergence rate bounds in more general cases.
3.2 Bernoulli Sampling with Constant Step Size: Second Set of Convergence Conditions
We consider algorithm (1) where $z_{t}$ is product of a fixed step size $\gamma$ and a Bernoulli random variable $\zeta_{t}$ with mean $\pi(x,y,\theta)$. Let $g_{t}=\zeta_{t}\nabla_{\theta}\ell(x_{t},y_{t},\theta_{t})$, which is random vector because $\zeta_{t}$ is a random variable and $(x_{t},y_{t})$ is a sampled point. Following Liu and Li (2023), we note that the algorithm (1) is a stochastic gradient descent algorithm with respect to an objective function $\tilde{\ell}$ with gradient
@@ 3.2 Bernoulli Sampling with Constant Step Size: Second Set of Convergence Conditions
where the expectation is with respect to $x$ and $y$. This observation allows us to derive convergence rate results by deploying convergence rate results for stochastic gradient algorithm known to hold under various assumptions on function $\tilde{\ell}$, variance of stochastic gradient vector and step size. A function $\tilde{\ell}$ satisfying condition (7) is referred to as equivalent loss in Liu and Li (2023).
@@
The function ~$\tilde{\ell}(\theta)$ inherits some properties of functions $\ell(x,y,\theta)$, for $(x,y)\in\mathcal{X}\times\mathcal{Y}$. For example, if $\ell(x,y,\theta)$ is a convex function, for every $(x,y)\in\mathcal{X}\times\mathcal{Y}$, then $\tilde{\ell}$ is a convex function.
The framework for establishing convergence rates allows us to accommodate different sampling strategies and loss functions. For example, it accommodates binary classification for binary cross-entropy loss function and sampling proportional to absolute error loss, as shown in the following.
For binary classification, we model the prediction probability of the positive label as $\sigma(x^{\top}\theta)$, where $\sigma:\mathbb{R}_{+}\rightarrow[0,1]$ is an increasing function, and $\sigma(-u)+\sigma(u)=1$ for all $u\in\mathbb{R}$. The absolute error loss takes value $1-\sigma(x^{\top}\theta)$ if $y=1$ or value $\sigma(x^{\top}\theta)$ if $y=-1$, which corresponds to $1-\sigma(yx^{\top}\theta)$. The binary cross-entropy loss for a point $(x,y)$ under model parameter $\theta$ can be written as $\ell(x,y,\theta)=-\log(\sigma(yx^{\top}\theta))$. Hence, absolute error loss-based sampling corresponds to the sampling probability function $\pi(x)=1-e^{-x}$.
The next lemma allows to derive convergence rate results for expected loss with respect to loss function $\ell$ by applying convergence rate results for expected loss with respect to loss function $\tilde{\ell}$ (which, recall, is the equivalent loss function).
@@
Assume that $\ell$ is a convex function, $\tilde{\ell}$ is $L$-smooth, $\Theta_{0}$ is a convex set, $S=\sup_{\theta\in\Theta_{0}\cap\Theta}||\theta-\theta_{1}||$, and $\mathbb{E}[\pi(\ell(x,\theta))||\nabla_{\theta}\ell(x,y,\theta)||^{2}]-||\nabla_{\theta}\tilde{\ell}(\theta)||^{2}\leq\sigma_{\pi}^{2}$. Then, algorithm (1) with $\gamma=1/(L+(\sigma/R)\sqrt{n/2})$,
$\mathbb{E}[\ell(\bar{\theta}_{n})]\leq\mathbb{E}\left[\frac{1}{n}\sum_{t=1}^{n}\ell(\theta_{t})\right]$
$\leq$ $\inf_{\theta}\Pi^{-1}(\tilde{\ell}(\theta))+\Pi^{-1}\left(\frac{\sqrt{2}S\sigma_{\pi}}{\sqrt{n}}\right)+\Pi^{-1}\left(\frac{LS^{2}}{n}\right).$
Note that the bound on the expected loss in Theorem 3.7 depends on $\pi$ through $\Pi^{-1}$ and $\sigma_{\pi}^{2}$. Obviously, we have a bound depending on $\pi$ only through $\Pi^{-1}$ by upper bounding $\sigma_{\pi}^{2}$ with $\sup_{\theta\in\Theta_{0}}\mathbb{E}[||\nabla_{\theta}\ell(x,y,\theta)||^{2}]$.
@@
$\mathbb{E}[\ell(\bar{\theta}_{n})]\leq\mathbb{E}\left[\frac{1}{n}\sum_{t=1}^{n}\ell(\theta_{t})\right]$
$\leq$ $\inf_{\theta}\Pi^{-1}(\tilde{\ell}(\theta))+2^{5/4}\sqrt{S\sigma_{\pi}}\frac{1}{\sqrt[4]{n}}+2\sqrt{L}S\frac{1}{\sqrt{n}}.$
@@
3.3 Bernoulli Sampling with Stochastic Polyak’s Step Size: A New Algorithm
In this section we propose a new algorithm that combines a Bernoulli sampling and an adaptive stochastic gradient descent update, and provide a convergence rate guarantee. The algorithm is defined as the stochastic gradient descent algorithm (1) with stochastic step size $z_{t}$ being a binary random variable that takes value $\gamma_{t}:=\zeta(x_{t},y_{t},\theta_{t})/\pi(x_{t},y_{t},\theta_{t})$ with probability $\pi(x_{t},y_{t},\theta_{t})$ and takes value 00 otherwise, where $\pi$ is some function and if $||\nabla_{\theta}\ell(x,y,\theta)||>0$,
@@ 3.3 Bernoulli Sampling with Stochastic Polyak’s Step Size: A New Algorithm
and, $\zeta(x,y,\theta)=0$, otherwise, for constants $\beta,\rho>0$ and
@@
The expected step size $\zeta(x,y,\theta)$ corresponds to the adaptive step size proposed by Polyak (Loizou et al., 2021). When $\pi(x,y,\theta)=1$ for all $x,y,\theta$, the algorithm reduces to Polyak’s adaptive step size algorithm (Loizou et al., 2021). Our algorithm introduces a sampling component and re-weighting of the update to ensure that the step size remains according to the stochastic Polyak’s step size in expectation. For many loss functions $\inf_{\theta^{\prime}}\ell(x,y,\theta)=0$, for every $x,y$. In these cases, $\psi(x,y,\theta)=||\nabla_{\theta}\ell(x,y,\theta)||^{2}/\ell(x,y,\theta)$. For instance, for binary cross-entropy loss function, $\inf_{\theta^{\prime}}\ell(x,y,\theta)=\inf_{\theta^{\prime}}(-\log(\sigma(yx^{\top}\theta^{\prime})))=0$, for all $x,y$.
We next show a convergence rate guarantee.
Lemma 3.10.
@@
$\mathbb{E}\left[\frac{1}{n}\sum_{t=1}^{n}(\ell(x_{t},y_{t},\theta_{t})-\ell(x_{t},y_{t},\theta^{*}_{t}))\right]$
$\leq$ $\frac{\rho\beta}{c\kappa}\Lambda^{*}+\frac{1}{2c\kappa}||\theta_{1}-\theta^{*}||^{2}\frac{1}{n}$
@@
By taking $\rho=1/(2L)$, the bound on the expected average loss boils down to $\Lambda^{*}/c+(L/(c\beta))||\theta_{1}-\theta^{*}||^{2}/n$.
Notably, under the condition on the sampling probability in Lemma 3.10, the convergence rate is of order $O(1/n)$. A similar bound is known to hold for stochastic gradient descent with adaptive stochastic Polyak step size for the finite-sum problem, as seen in Theorem 3.4 of Loizou et al. (2021). Lemma 3.10 applies to stream-based stochastic gradient descent algorithms and, importantly, allows for Bernoulli sampling with a sampling probability function $\pi$.
We next show instances of $\pi$ that satisfy condition (10).
Constant Sampling Probability
Assume that $\kappa\leq 2(1-c)$ and sampling probability function $\pi$ is constant with value at least $\frac{\kappa}{2(1-c)}$. Then, the expected loss bound in Lemma 3.10 holds.
Power Function of the Expected Step Size
Consider $\pi$ defined as
$\pi(x,y,\theta)=\zeta(x,y,\theta)^{\eta}$ (11)
where $\eta\geq 0$ and $\rho\beta\in(0,1]$. This sampling probability function is an increasing function of $1/\psi(x,y,\theta)$, i.e. $(\ell(x,y,\theta)-\inf_{\theta^{\prime}}\ell(x,y,\theta))/||\nabla_{\theta}\ell(x,y,\theta)||^{2}$. In other words, the sampling probability is an increasing function of the loss normalized by the squared norm of the gradient. Assume that
$\beta^{1-\eta}\leq 2(1-c)\left(\frac{1}{2L}\right)^{\eta}.$ (12)
Then, the loss bound in Theorem 3.10 holds.
Linear Classifiers: Binary Classification
For linear classifiers, if $\ell(u)$ is an $L^{\prime}$-smooth function in $u$, then $\ell(yx^{\top}\theta)$ is an $L^{\prime}||x||^{2}$-smooth function in $\theta$. To satisfy condition (10), it suffices that
$\pi(x,y,\theta)\geq\frac{\beta}{2(1-c)}\min\left\{\rho R^{2}h(yx^{\top}\theta),1\right\}$
where $h(u):=\ell^{\prime}(u)^{2}/\ell(u)$ and note that $||\nabla_{\theta}\ell(x,y,\theta)||^{2}/\ell(x,y,\theta)=h(yx^{\top}\theta)||x||^{2}$. By setting $\rho=1/(2L)$ with $L=L^{\prime}R^{2}$, we get $\rho R^{2}=1/(2L^{\prime})$.
For logistic regression and the binary cross-entropy loss function, the function $h(u)$ can be written as follows
$h(u)=\frac{1}{(1+e^{u})^{2}\log(1+e^{-u})}$ (13)
which is increasing in $u$ for $u\leq 0$ and decreasing in $u$ otherwise. Note that $h(u)=(1-e^{-\ell(u)})^{2}/\ell(u)$. Further properties of function $h$, deferred to Appendix A.13, establish two key facts: (a) Condition of Lemma 3.10 is satisfied with $\rho=1/(2L)$ by sampling proportional to absolute error loss $\pi^{*}(u)=\omega(1-\sigma(u))$ with $\beta/(4(1-c)L^{\prime})\leq\omega\leq 1$, and (b) Condition of Lemma 3.10 is satisfied by uncertainty sampling according to, for any fixed $a\in(0,1/2]$,
$\pi^{*}(u)=\frac{\beta}{2(1-c)}\min\left\{\rho R^{2}\frac{1}{H(a)+(1-a)|u|},1\right\}$
where $H(a)=a\log(1/a)+(1-a)\log(1/(1-a))$.
Linear Classifiers: Multi-class Classification
Condition (10) holds for some loss and uncertainty-based sampling methods. Refer to Appendix A.15 for detailed discussion.
@@ 4 Numerical Results
Figure 2: Average cross-entropy progressive loss of Polyak’s step size Loizou et al. (2021) compared to SGD with constant step size, for 1% and 10% sampling from the mushrooms data.
@@
Figure 3: Convergence in terms of average cross-entropy progressive loss of random sampling, loss-based sampling based on the absolute error loss, and our proposed adaptive step size variant of absolute error loss sampling.
@@
We adopt the setting from Loizou et al. (2021) and utilize the mushroom classification dataset (Chang and Lin, 2011) for binary classification tasks using RBF kernels and logistic regression. This dataset comprises $n=6499$ instances and $d=6499$ dimensions. In our training approach, we deliberately confine the training process to a single epoch. Throughout this epoch, we sequentially process each data instance, compute the loss for each individual instance, and subsequently update the model’s weights. This approach, known as progressive validation (Blum et al., 1999), enables us to monitor the evolution of the average loss. The constraint to a single epoch ensures that we calculate losses only for instances that haven’t influenced model weights. For each sampling scheme, we conduct a hyper-parameter sweep to minimize the average progressive loss.
Loizou et al. (2021) demonstrated that stochastic gradient descent with a step size corresponding to their stochastic Polyak’s step size (SPS) converges faster than gradient descent. Figure 2 illustrates that these findings extend to scenarios where we selectively sample from the dataset rather than training on the full dataset, and the step size is according to stochastic Polyak’s step size only in expectation.  [^1] We will make the code and the dataset available for download in the final version of the manuscript.
Figure 3 demonstrates that our proposed active learning algorithm, with adaptive Polyak step size according to SPS and sampling according to $\pi(x)=1-e^{-x}$, leads to faster convergence than the traditional loss-based sampling. The figure also shows that the traditional loss-based sampling approach converges more rapidly than random sampling.
We also simulate a noisy estimator of the absolute error loss for the proposed loss-based sampling with adaptive Polyak’s step size. This corresponds to the loss-based active learning setting in practice, where the loss is not known prior to label acquisition and can only be estimated. We model an unbiased noisy estimator ^$\hat{\ell}_{\mathrm{abs}}$ of the absolute error loss $\ell_{\mathrm{abs}}\in[0,1]$ as a random variable following the beta distribution, denoted as ^$\hat{\ell}_{\mathrm{abs}}\sim\mathrm{Beta}(\alpha,\beta)$, where $\alpha$ and $\beta$ are parameters set to ensure $\mathbb{E}[\hat{\ell}_{\mathrm{abs}}]=\ell_{\mathrm{abs}}$. The variance of the noise can be controlled by the tuning parameter $\alpha$, given by $\mathrm{var}[\hat{\ell}_{\mathrm{abs}}]=\frac{\ell_{\mathrm{abs}}(1-\ell_{\mathrm{abs}})}{\alpha+\ell_{\mathrm{abs}}}$. Figure 4). Figure 4 shows that the convergence results remain robust against noise in estimating the absolute error loss for a wide range of values for $\alpha\geq 1$.

### v3

@@ 3.1 Constant-Weight Sampling
Theorem 3.1, and our other results for linear classifiers and linearly separable datasets, are established using a convergence rate lemma, which is presented in Appendix A.2, along with its proof. This lemma generalizes the conditions used to establish the convergence rate for an uncertainty-based sampling algorithm by Raj and Bach (2022), with the sampling probability function $\pi(u)=1/(1+\mu|u|)$, for some constant $\mu>0$. It can be readily shown that Theorem 3.1 in Raj and Bach (2022) follows from our convergence rate lemma with the training loss function corresponding to the squared hinge loss function and the evaluation loss function (used for convergence rate guarantee) corresponding to the hinge loss function. Further details on the convergence rate lemma are discussed in Appendix A.2.1.
The convergence rate conditions for multi-class classification with the set of classes $\mathcal{Y}$ are the same as for binary classification, with $u(x,y,\theta):=x^{\top}\theta_{y}-\max_{y^{\prime}\in\mathcal{Y}\setminus\{y\}}x^{\top}\theta_{y^{\prime}}$, except for an additional factor of $2$ in one of the conditions (see Lemma A.10 in the Appendix). Hence, all the observations remain valid for the multi-class classification case.
More General Classifiers and Datasets
We consider algorithm (1) where $z_{t}$ is product of a fixed step size $\gamma$ and a Bernoulli random variable $\zeta_{t}$ with mean $\pi(x,y,\theta)$. Let $g_{t}=\zeta_{t}\nabla_{\theta}\ell(x_{t},y_{t},\theta_{t})$, which is random vector because $\zeta_{t}$ is a random variable and $(x_{t},y_{t})$ is a sampled point. Following Liu and Li (2023), we note that the algorithm (1) is an SGD algorithm with respect to an objective function $\tilde{\ell}$ with gradient
@@
where the expectation is with respect to $x$ and $y$. This observation allows us to derive convergence rate results by deploying convergence rate results that are known to hold for SGD under various assumptions on function $\tilde{\ell}$, variance of stochastic gradient vector and step size. A function $\tilde{\ell}$ satisfying condition (3) is referred to as an equivalent loss in Liu and Li (2023).
@@
If $\ell(x,y,\theta)$ is a convex function, for every $(x,y)\in\mathcal{X}\times\mathcal{Y}$, then $\tilde{\ell}$ is a convex function.
This framework for establishing convergence rates allows us to accommodate different sampling strategies and loss functions. The next lemma allows us to derive convergence rate results for expected loss with respect to loss function $\ell$ by applying convergence rate results for expected loss with respect to loss function $\tilde{\ell}$ (which, recall, is the equivalent loss function).
@@
Assume that $\ell$ is a convex function, $\tilde{\ell}$ is $L$-smooth, $\Theta_{0}$ is a convex set, $S=\sup_{\theta\in\Theta_{0}}||\theta-\theta_{1}||$, and $\mathbb{E}[\pi(\ell(x,\theta))||\nabla_{\theta}\ell(x,y,\theta)||^{2}]-||\nabla_{\theta}\tilde{\ell}(\theta)||^{2}\leq\sigma_{\pi}^{2}$. Then, for algorithm (1) with $\gamma=1/(L+(\sigma/R)\sqrt{n/2})$,
$\mathbb{E}[\ell(\bar{\theta}_{n})]\leq\mathbb{E}\left[\frac{1}{n}\sum_{t=1}^{n}\ell(\theta_{t})\right]\leq\inf_{\theta}\Pi^{-1}(\tilde{\ell}(\theta))+\Pi^{-1}\left(\frac{\sqrt{2}S\sigma_{\pi}}{\sqrt{n}}\right)+\Pi^{-1}\left(\frac{LS^{2}}{n}\right).$
Note that the bound on the expected loss in Theorem 3.3 depends on $\pi$ through $\Pi^{-1}$ and $\sigma_{\pi}^{2}$. Specifically, we have a bound depending on $\pi$ only through $\Pi^{-1}$ by upper bounding $\sigma_{\pi}^{2}$ with $\sup_{\theta\in\Theta_{0}}\mathbb{E}[||\nabla_{\theta}\ell(x,y,\theta)||^{2}]$.
@@
$\mathbb{E}[\ell(\bar{\theta}_{n})]\leq\mathbb{E}\left[\frac{1}{n}\sum_{t=1}^{n}\ell(\theta_{t})\right]\leq\inf_{\theta}\Pi^{-1}(\tilde{\ell}(\theta))+2^{5/4}\sqrt{S\sigma_{\pi}}\frac{1}{\sqrt[4]{n}}+2\sqrt{L}S\frac{1}{\sqrt{n}}.$
@@
3.2 Adaptive-Weight Sampling
In this section we propose the Adaptive-Weight Sampling (AWS) algorithm that combines Bernoulli sampling and an adaptive SGD update, and provide a convergence rate guarantee. The algorithm is defined by (1) with the stochastic step size $z_{t}$ being a binary random variable that takes value $\gamma_{t}:=\zeta(x_{t},y_{t},\theta_{t})/\pi(x_{t},y_{t},\theta_{t})$ with probability $\pi(x_{t},y_{t},\theta_{t})$ and takes value 00 otherwise, where $\pi$ is some sampling probability function. Here, $\zeta(x,y,\theta)$ is the expected SGD (1) step size, defined as
@@ 3.2 Adaptive-Weight Sampling
whenever $||\nabla_{\theta}\ell(x,y,\theta)||>0$ and $\zeta(x,y,\theta)=0$ otherwise, for constants $\beta,\rho>0$, where
@@
The expected step size $\zeta(x,y,\theta)$ corresponds to the stochastic Polyak’s step size used by a gradient descent algorithm proposed by (Loizou et al., 2021), which is accommodated as a special case when $\pi(x,y,\theta)=1$ for all $x,y,\theta$. AWS introduces a sampling component and re-weighting of the update to ensure that the step size remains according to the stochastic Polyak’s step size in expectation. For many loss functions, $\inf_{\theta^{\prime}}\ell(x,y,\theta)=0$, for every $x,y$. In these cases, $\psi(x,y,\theta)=||\nabla_{\theta}\ell(x,y,\theta)||^{2}/\ell(x,y,\theta)$. For instance, for binary cross-entropy loss function, $\inf_{\theta^{\prime}}\ell(x,y,\theta)=\inf_{\theta^{\prime}}(-\log(\sigma(yx^{\top}\theta^{\prime})))=0$, for all $x,y$.
We next show a convergence rate guarantee for AWS.
Theorem 3.6.
@@
$\mathbb{E}\left[\frac{1}{n}\sum_{t=1}^{n}(\ell(x_{t},y_{t},\theta_{t})-\ell(x_{t},y_{t},\theta^{*}_{t}))\right]\leq\frac{\rho\beta}{c\kappa}\Lambda^{*}+\frac{1}{2c\kappa}||\theta_{1}-\theta^{*}||^{2}\frac{1}{n}$
@@
The bound on the expected average loss in Theorem 3.6 boils down to $\Lambda^{*}/c+(L/(c\beta))||\theta_{1}-\theta^{*}||^{2}/n$ by taking $\rho=1/(2L)$. Notably, under the condition on the sampling probability in Theorem 3.6, the convergence rate is of order $O(1/n)$. A similar bound is known to hold for SGD with adaptive stochastic Polyak step size for the finite-sum problem, as seen in Theorem 3.4 of Loizou et al. (2021). A difference is that Theorem 3.6 allows for sampling of the points.
Loss and Uncertainty-based Sampling for Linear Binary Classifiers
We consider linear binary classifiers, focusing particularly on logistic regression and the binary cross-entropy training loss function. The following corollaries of Theorem 3.6 hold for sampling proportional to absolute error loss and an uncertainty-based sampling probability function, respectively.
Corollary 3.7.
For sampling proportional to absolute error loss, $\pi(u)=\omega(1-\sigma(u))$, with $\beta/(4(1-c)L^{\prime})\leq\omega\leq 1$ and $\rho=1/(2L)$, the bound on the expected loss in Theorem 3.6 holds.
Corollary 3.8.
For the uncertainty-based sampling according to
$\pi(u)=\frac{\beta}{2(1-c)}\min\left\{\rho R^{2}\frac{1}{H(a)+(1-a)|u|},1\right\}$
where $a\in(0,1/2]$ and $H(a)=a\log(1/a)+(1-a)\log(1/(1-a))$, the bound on the expected loss in Theorem 3.6 holds.
Other Cases
For a constant sampling probability function with a value of at least $\kappa/(2(1-c))$, condition (6) holds when $\kappa\leq 2(1-c)$. When $\pi(x,y,\theta)=\zeta(x,y,\theta)^{\eta}$, where $\eta\geq 0$ and $\rho\beta\in(0,1]$, condition (6) holds under $\beta^{1-\eta}\leq 2(1-c)(1/(2L))^{\eta}$, as shown in Appendix A.14. Condition (6) also holds for an uncertainty-based sampling in multi-class classification, as shown in Appendix A.15.
@@ 4 Numerical Results
(a) Mushrooms dataset
(b) MNIST classification of 3s vs 5s
(c) Parkinsons dataset
(d) Splice dataset
(e) Tictactoe dataset
(f) Credit dataset
@@
Figure 1: Convergence in terms of average cross-entropy progressive loss of random sampling, loss-based sampling based on the absolute error loss, and our proposed algorithm (loss-based sampling with stochastic Polyak’s step size). Our proposed algorithm outperforms the baselines in most cases.
In this section we evaluate our AWS algorithm, defined in Section 3.2. In particular, we focus on an instance of AWS with stochastic Polyak’s expected step size for logistic regression and the loss-based sampling proportional to absolute error loss, which we refer to as Adaptive-Weight Sampling - Polyak Absloss (AWS-PA). By, Corollary 3.7, AWS-PA converges according to Theorem 3.6. Here we demonstrate convergence on real-world datasets and compare with other algorithms.
We use a modified version of the mushroom binary classification dataset (Chang and Lin, 2011) that was used by Loizou et al. (2021) for evaluation of their algorithm. This modification uses RBF kernel features, resulting in a linearly separable dataset for a linear classifier like logistic regression. Furthermore, we include five datasets that we selected at random from the 44 real-world datasets that were used in Yang and Loog (2018), a benchmark study of active learning for logistic regression: MNIST 3 vs 5 LeCun et al. (1998), parkinsons Little et al. (2007), splice Noordewier et al. (1990), tictactoe Aha (1991), and credit Quinlan (1987). While these datasets are not necessarily linearly separable, Yang and Loog (2018) has shown that logistic regression achieves a good quality-of-fit.
@@
In our evaluation, we deliberately confine the training to a single epoch. Throughout this epoch, we sequentially process each data instance, compute the loss for each individual instance, and subsequently update the model’s weights. This approach, known as progressive validation (Blum et al., 1999), enables us to monitor the evolution of the average loss. The constraint to a single epoch ensures that we calculate losses only for instances that haven’t influenced model weights. For each sampling scheme, we conduct a hyper-parameter sweep to minimize the average progressive loss and apply a procedure to ensure that all algorithms sample comparable numbers of instances. In Appendix B.1 we include further details on this procedure, the hyper-parameter tuning, and other aspects of the experimental setup.
Figure 1 demonstrates that AWS-PA leads to faster convergence than the traditional loss-based sampling with a constant step size (akin to Yoo and Kweon (2019)). It also shows that the traditional loss-based sampling approach converges more rapidly than random sampling on five of the six datasets. These results are obtained under a hyper-parameter tuning such that different algorithms have comparable data sampling rates. We provide additional experimental results that demonstrate the efficiency of AWS-PA in Appendix B.2.
(a) Mushrooms dataset
(b) MNIST classification of 3s vs 5s
(c) Parkinsons dataset
(d) Splice dataset
(e) Tictactoe dataset
(f) Credit dataset
Figure 2: Active learning sampling based on an estimator of the absolute error loss performs on par with the sampling based on the ground truth value of absolute error loss.
In active learning applications, the true loss of a point cannot be computed before the corresponding label is obtained. Hence, in practice we do not know the true loss at the moment of making the sampling decision. Therefore, we assess the effect of using a loss estimator, instead of using the true loss values. We use a Random Forest regressor to estimate absolute error loss based on the same set of features as the target model and the target’s model prediction as an extra feature. We retrain this estimator on every sampling step using the labeled points observed so far. Figure 2 demonstrates that AWS-PA with the estimated absolute error losses performs similarly on all datasets to AWS-PA with the true absolute error losses. Moreover, for a majority of the datasets, the two variants of AWS-PA achieve similar data sampling rates; this is shown Appendix B.3 along with further discussion.